# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 3.2. Application to Ground State Calculations in Quantum Chemistry
A fundamental problem in computational quantum chemistry is the numerical approximation of the lowest eigenvalue (or so-called ground state energy) of the electronic Hamiltonian– a self-adjoint operator acting on an $L^{2}$-type Hilbert space of antisymmetric functions. The aim of this section is to briefly describe how this energy minimization problem can be reformulated to fit the framework of tensor train approximations described in Sections 2 using the so-called second quantization formalism. Subsequently, in the forthcoming Section 4, we will demonstrate the performance of the A2DMRG Algorithm 2 for the resolution of this energy minimization problem. For simplicity, we will do so without referring to the spin variable but we emphasize that all numerical tests in Section 4 do, in fact, include the spin variable.
We begin by considering a so-called single-particle basis, i.e., a finite collection of $L^{2}(\mathbb{R}^{3})$-orthonormal basis functions $\{\phi_{\mu}\}_{\mu=1}^{d}\subset H^{1}(\mathbb{R}^{3})$. In quantum mechanics, a fundamental distinction is made between so-called bosonic and fermionic particles, the latter obeying the so-called Pauli-exclusion principle and thus being described in terms of antisymmetric functions. Since electrons are fermionic particles, we next introduce, for each $\ell\in\{1,\ldots,d\}$, the set $\mathcal{B}_{\ell}$ of antisymmetric functions defined as
@@
We also define the finite-dimensional vector space $\mathcal{V}_{\ell}$ as
[…]
Here, the symbol $\wedge$ is used to denote the antisymmetric tensor product (see, e.g., [35, Chapter One]) so that, in the present context, we have
$H^{1}(\mathbb{R}^{3\ell})\ni$ $\Phi=\phi_{\mu_{1}}\wedge\phi_{\mu_{2}}\ldots\wedge\ldots\phi_{\mu_{\ell}}$
$\iff$ $\Phi(\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{\ell})=\frac{1}{\sqrt{\ell!}}\sum_{\pi\in{\mathrm{S}}(\ell)}(-1)^{\rm sgn(\pi)}\otimes_{i=1}^{\ell}\phi_{mu_{i}}\big{(}\mathbb{x}_{\pi(i)}\big{)},$
where ${\mathrm{S}}(\ell)$ is the permutation group of order $\ell$, and ${\rm sgn(\pi)}$ denotes the signature of $\pi~{}\in~{}{\mathrm{S}}(\ell)$.
In the quantum chemistry literature, the set $\mathcal{B}_{\ell}$ is typically referred to as an $\ell$-particle basis set while the vector space $\mathcal{V}_{\ell}$ is the corresponding $\ell$-particle function space. Equipped with these sets and spaces, we now introduce the so-called Fermionic Fock space as the $2^{d}$-dimensional vector space defined as
(20) $\mathcal{F}_{d}:=\bigoplus_{\ell=0}^{d}\mathcal{V}_{\ell},\qquad\text{equipped with the natural $\bigoplus_{\ell=0}^{d}L^{2}(\mathbb{R}^{3\ell})$ norm.}$.
Of particular importance are the so-called fermionic creation and annihilation operators acting on the Fock space $\mathcal{F}_{d}$ (see, e.g, [16, Chapter 1] for further details). For any $\ell\in\{1,\ldots,d\}$, the $\ell^{\rm th}$ fermionic creation operator $a_{\ell}^{\dagger}\colon\mathcal{F}_{d}\rightarrow\mathcal{F}_{d}$ is defined as
@@
[…]
@@
$\mathcal{H}$ $=\sum_{i,j=1}^{d}{h}_{ij}a_{i}^{\dagger}a_{j}+\sum_{i,j,k,\ell=1}^{d}{V}_{ijk\ell}a_{i}^{\dagger}a_{j}^{\dagger}a_{k}a_{\ell},$
where
[…]
Assuming that the molecular system is composed precisely of $N$ electrons, the so-called ground-state problem for this given molecular system consists of finding the lowest eigenvalue of the second quantized Hamiltonian $\mathcal{H}$, subject to the constraint that the corresponding eigenvector describes an $N$-electron system. This constraint can be imposed by requiring that the sought-after energy minimizer be an eigenfunction of the so-called particle number operator $\mathcal{N}=\sum_{j=1}^{d}a_{j}^{\dagger}a_{j}\colon\mathcal{F}_{d}\rightarrow\mathcal{F}_{d}$ with eigenvalue precisely $N$. Consequently, the ground state problem can therefore be expressed as a constrained minimization problem of the form
(21) $\mathcal{E}_{\rm GS}=\min_{0\neq\Phi\in\mathcal{F}_{d}}\frac{\langle\Phi,\mathcal{H}\Phi\rangle_{\mathcal{F}_{d}}}{\langle\Phi,\Phi\rangle_{\mathcal{F}_{d}}}\qquad\text{subject to the constraint }\mathcal{N}\Phi:=\sum_{j=1}^{d}a_{j}^{\dagger}a_{j}\Phi=N\Phi.$
@@
Consequently, the minimization problem (21), which is a priori posed on the Fock space $\mathcal{F}_{d}$ can equivalently be formulated over the space $\bigotimes_{j=1}^{d}\mathbb{R}^{2}$ of order-$d$ tensors:
@@
The minimization problem (22) can now be approximated using the tensor train format and alternating minimization algorithms described in detail in Sections 2 and 3. Note, however, that these sections did not include any discussion on the imposition of the particle number constraint included in the minimization problem (21). Let us, therefore, briefly describe how this additional constraint can be dealt with in the tensor train format.
It has been shown by Bachmayr, Götte, and Pfeffer [1] that eigenfunctions of the particle number operator $\iota\circ\mathcal{N}\circ\iota^{-1}\rightarrow\mathbb{R}^{2^{d}}\times\mathbb{R}^{2^{d}}$ are characterized by a block-sparse structure in their tensor-train decomposition. Therefore, to approximate the minimization problem (22) on a tensor manifold $\mathcal{T}_{\mathbb{r}},\mathbb{r}\in\mathbb{N}^{d-1}$ using, e.g., classical one-site DMRG, we first initialize the algorithm with some $\mathbb{U}^{(0)}\in\mathcal{U}_{\mathbb{r}}$ that possess the requisite block-sparse structure. This guarantees that $\tau(\mathbb{U}^{(0)})$ is an eigenfunction of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$. Next, assume that we are performing micro-iteration $k+1$ of the one-site DMRG algorithm, and that the previous iterate $\mathbb{U}^{(k)}\in\mathcal{U}_{\mathbb{r}}$ is $k+1$-orthogonal and possesses the requisite block-sparse structure. As an alternative to Lemma 10, we can characterize the $k+1$ DMRG micro-iterate $\mathbb{U}^{(k+1)}\in\mathcal{U}_{\mathbb{r}}$ through projection operators on the tangent space $\mathbb{T}_{\mathbb{U}^{(k)}}\mathcal{T}_{\mathbb{r}}$.
To this end, for each $j\in\{1,\ldots,d\}$, let us denote by $\mathbb{U}^{(k)}_{\leq j}\in\mathbb{R}^{n_{1}\ldots n_{i}\times r_{i}}$ and $\mathbb{U}^{(k)}_{\geq j}\in\mathbb{R}^{r_{j-1}\times n_{j}\ldots n_{d}}$ the matrices obtained by contracting and reshaping the first $j$ TT cores and the last $d-j+1$ TT cores, respectively, of $\mathbb{U}^{(k)}$. We can then introduce, for each $j\in\{1,\ldots,d\}$ the projection operators $\mathbb{P}_{\leq j},\mathbb{P}_{\geq j},\mathbb{P}_{j}\colon\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{T}_{\mathbb{U}^{(k)}}\mathcal{T}_{\mathbb{r}}$ as
@@
Here, $\text{\rm Ten}_{j}$ is a reshape operation that takes a matrix in $\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{d}}$ and reshapes it into a tensorial element of $\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$, and we adopt the convention that $\mathbb{P}_{\leq 0}=\mathbb{P}_{\geq d+1}=1$.
Recalling Lemma 10, it is now possible to show (see, e.g., [1]) that
$\tau(\mathbb{U}^{(k+1)})\text{ is the lowest eigenfunction of }\mathbb{P}_{j}^{*}\circ\iota\circ\mathcal{H}\circ\iota^{-1}\circ\mathbb{P}_{j}.$
@@
Finally, let us turn to the question of particle-number preservation in the A2DMRG Algorithm 2. As in the classical DMRG algorithms, we begin with an initialization $\mathbb{U}^{(0)}\in\mathcal{U}_{\mathbb{r}}$ that possesses the requisite block-sparse structure which guarantees that $\tau(\mathbb{U}^{(0)})$ is an eigenfunction of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$. We then perform $d$ (resp. $d-1$ in the two-site case) parallel DMRG micro-iterations resulting in updated tensors $\{\tau(\mathbb{V}^{(j)}\}_{j=1}^{d}$ (resp. $\{\tau(\mathbb{W}^{(j,j+1})\}_{j=1}^{d-1}$). As argued above, each DMRG micro-iteration with an appropriately orthogonalized and particle-preserving initialization results in a particle-preserving tensor. Therefore, all tensors $\{\tau(\mathbb{V}^{(j)}\}_{j=1}^{d}$ (resp. $\{\tau(\mathbb{W}^{(j,j+1)}\}_{j=1}^{d-1}$) are eigenfunctions of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$ and so are linear combinations of such tensors. It therefore suffices to consider the effect of the compression step (Step 4 in Algorithm 2) which projects a linear combination $\sum_{j=1}^{d}c_{j}^{*}\tau(\mathbb{V}^{(j)},~{}\{c_{j}^{*}\}_{j=1}^{d}\subset\mathbb{R}$ (resp. $\sum_{j=1}^{d-1}d_{j}^{*}\tau(\mathbb{W}^{(j,j+1)},~{}\{d_{j}^{*}\}_{j=1}^{d-1}\subset\mathbb{R}$) onto some low-rank tensor manifold $\mathcal{T}_{\mathbb{r}}$. Such a compression step can be performed using either the particle-preserving version of the TT-rounding algorithm [1, Algorithm 3], or more-generally, by solving a best-approximation problem using standard tensor train algorithms (such as alternating least squares), which are known to be particle-preserving [1, Section 6].
@@ 4. Numerical Experiments
For the purpose of the subsequent numerical study, we consider the tensor formulation of the ground state energy minimization problem (22) in the case of several so-called strongly correlated molecular systems. This includes linear chains of hydrogen atoms, i.e., molecular systems of the form H2p for $p\in\{3,4,5,6\}$ as well as the C2 and N2 dimers. The bond lengths in each Hydrogen chain and the C2 dimer are set to one Angstrom while the bond length for the N2 dimer is set to $1.0976$ Angstrom. For each molecular system, the $L^{2}(\mathbb{R}^{3})$-orthonormal single-particle basis $\{\phi_{j}\}_{j=1}^{d}$ is obtained by diagonalizing the Hartree-Fock (or Mean-Field) operator in the minimal STO-3G atomic basis. This implies, in particular, the relation $d=4p$ for each $p\in\{3,\ldots,6\}$ and $d=20$ for C2 and N2 so that our numerical experiments involve tensors of minimal order $d=12$ in the case of $H_{6}$ and maximal order $d=24$ in the case of $H_{12}$. The input data, i.e., the second quantized electronic Hamiltonian, and the reference minimal energies are obtained from the PySCF Python module for quantum chemistry [40] while the DMRG simulations themselves are run using an in-house Julia module developed and described in [2].
Figure 1. Convergence plot of the classical two-site DMRG and two-site A2DMRG algorithms for a linear, stretched H6 and H8 chain. The tolerances for the eigenvalue solvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive sweeps/global iterations was smaller than $10^{-6}$.
@@
Figure 2. Convergence plot of the classical two-site DMRG and two-site A2DMRG algorithms for a linear, stretched H10 and H12 chain. The tolerances for the eigenvalue solvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive sweeps/global iterations was smaller than $10^{-6}$.
Figures 1-3 display the resulting error plots as a function of either the number of global iterations or the cost per processor. The term global iteration refers to one sweep of the classical DMRG algorithm or one iteration of the A2DMRG Algorithm 2, both of which consists of solving exactly $d-1$ local minimization problems. The cost per processor is computed by evaluating, through the course of each global iteration, the number of tensor contractions, matrix-vector products, and linear algebra operations such as QR and singular value decompositions. In the case of the classical DMRG algorithm, which is sequential in nature, we assume that all operations are performed on a single processor. In the case of the A2DMRG Algorithm 2, we assume that Steps 2 and 3, i.e., the DMRG micro-iterations and the solution of the second-level minimization problems are computed as much as possible in parallel. In other words, we assume that each of the $(d-1)$ DMRG micro-iterations in Step 2 of Algorithm 2 is performed on an independent processor, and each entry of the $d\times d$ symmetric coarse matrices appearing in the second-level minimization problem (Recall Equation 18) is computed on an independent processor. The cost per processor is then computed by taking the processor with the highest cost for Steps 2 and 3, and adding to it the cost of Steps 1 and 4 of Algorithm 2, the latter two steps being sequential in nature.
Figure 3. Convergence plot of the classical DMRG and A2DMRG algorithms for the C2 and N2 dimers. The tolerances for the eigenvalue solvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive sweeps/global iterations was smaller than $10^{-6}$.
A number of observations can now be made. First we note that the A2DMRG Algorithm 2 demonstrates slower convergence than the classical DMRG algorithm 1 as a function of the global iterations. Although, a rigorous quantification of the relative convergence rates of these two algorithms seems out of reach at the moment (indeed, even for the one-site DMRG algorithm, the local convergence analysis in [36] does not yield convergence rates), we feel that there is an intuitive explanation for the observed behavior: First, due to the sequential nature of the classical DMRG micro-iterations, the TT ranks of the ansatz grow very quickly and reach the maximal rank parameter $\mathbb{r}$ within two sweeps. On the other hand, the A2DMRG algorithm requires several global iterations before a maximal rank parameter is achieved thereby slowing the convergence rate. Second, it is well-known in the domain decomposition literature that multiplicative Schwarz methods (which can be seen as generalizations of the Gauss-Seidel method) converge, in general, faster than additive Schwarz methods (which can be seen as generalizations of the Jacobi method). Intuitively, therefore, the classical DMRG algorithm which can be viewed as a non-linear, overlapping multiplicative Schwarz method should also converge faster than the A2DMRG Algorithm 2 which can be interpreted as a non-linear, overlapping additive Schwarz method. Interestingly, such a domain-decomposition-based interpretation of DMRG has– to our knowledge– not been studied in the literature.
@@
Second, in spite of the previous observation on its relative convergence rate, we nevertheless observe that the A2DMRG Algorithm 2 exhibits a significant parallel speedup in terms of the cost per processor. The exact speed-up varies depending on the choice of ranks and molecule (the latter determines the order $d$ of the tensor), but ranges from a factor two to a factor eight. Part of the reason for this speed-up (which is lower than the theoretical optimal speed-up of factor $d/4$) is that the classical and additive DMRG algorithms seem to require a different number of Lanczos iterations for the solution of the micro-iterations. To explore this, we plot in Figures 4-6, the number of Lanczos iterations required for the DMRG micro-iterations over the course of the DMRG sweeps/global iterations. These results indicate that, in general, only a minority of local problems in the classical DMRG algorithm require more than a single Lanczos iteration. This is not the case for the A2DMRG algorithm, which typically requires several Lanczos iterations for each micro-iteration, thereby producing a non-optimal speed-up. However, we emphasize that the number of Lanczos iterations for the A2DMRG algorithm displays minimal growth as a function of the maximal TT ranks $\mathbb{r}$ and the order $d$ of the tensor.
A The classical two-site DMRG algorithm.
B The two-site A2DMRG algorithm.
C The classical two-site DMRG algorithm.
D The two-site A2DMRG algorithm.
A The classical two-site DMRG algorithm.
B The two-site A2DMRG algorithm.
C The classical two-site DMRG algorithm.
D The two-site A2DMRG algorithm.
Figure 6. The number of Lanczos iterations required for the micro-iterations in the classical DMRG and A2DMRG algorithms. All tolerances were set to $10^{-6}$.

### v2

@@ 3.2. Application to Ground State Calculations in Quantum Chemistry
A fundamental problem in computational quantum chemistry is the numerical approximation of the lowest eigenvalue (or so-called ground state energy) of the electronic Hamiltonian– a self-adjoint operator acting on an $L^{2}$-type Hilbert space of antisymmetric functions. In this section, we briefly recall how this energy minimization problem can be reformulated using the second quantization formalism to fit the framework of tensor train approximations described in Section 2. Subsequently, in Section 4, we will demonstrate the performance of the A2DMRG Algorithm 2 for the resolution of this energy minimization problem. For simplicity, our exposition ignores the spin variable but all numerical tests in Section 4 do, in fact, include spin.
We begin by considering a finite collection of $L^{2}(\mathbb{R}^{3})$-orthonormal basis functions $\{\phi_{\mu}\}_{\mu=1}^{d}\subset H^{1}(\mathbb{R}^{3})$ and we introduce, for each $\ell\in\{1,\ldots,d\}$, the set $\mathcal{B}_{\ell}$ of antisymmetric functions defined as
@@
Here, $\wedge$ denotes the antisymmetric tensor product (see, e.g., [37, Chapter 1]) so that
$\Phi=\phi_{\mu_{1}}\wedge\ldots\wedge\phi_{\mu_{\ell}}\iff\Phi(\mathbb{x}_{1},\ldots,\mathbb{x}_{\ell})=\frac{1}{\sqrt{\ell!}}\sum_{\pi\in{\mathrm{S}}(\ell)}(-1)^{\rm sgn(\pi)}\overset{\ell}{\underset{i=1}{\otimes}}\phi_{\mu_{i}}(\mathbb{x}_{\pi(i)}),$
with ${\mathrm{S}}(\ell)$ the permutation group of order $\ell$ and $\rm sgn(\pi)$} the signature of $\pi~\in~{\mathrm{S}}(\ell)$.
Denoting now $\mathcal{V}_{\ell}:=\text{\rm span}\mathcal{B}_{\ell}$ for all $\ell\in\{1,\ldots,d\}$ and setting $\mathcal{V}_{0}=\mathbb{R}$ by convention, we can introduce the so-called Fermionic Fock space as the $2^{d}$-dimensional Hilbert space given by
(20) $\mathcal{F}_{d}:=\bigoplus_{\ell=0}^{d}\mathcal{V}_{\ell},\qquad\text{equipped with the natural ~$\bigoplus_{\ell=0}^{d}L^{2}(\mathbb{R}^{3\ell})$ norm.}$
Of particular importance are the so-called fermionic creation and annihilation operators acting on the Fock space $\mathcal{F}_{d}$ (see, e.g, [18, Chapter 1] for further details). For any $\ell\in\{1,\ldots,d\}$, the fermionic creation operator $a_{\ell}^{\dagger}\colon\mathcal{F}_{d}\rightarrow\mathcal{F}_{d}$ is defined as
@@
The fermionic annihilation operator $a_{\ell}\colon\mathcal{F}_{d}\rightarrow\mathcal{F}_{d}$ is then the $\overset{d}{\underset{\ell=0}{\bigoplus}}L^{2}(\mathbb{R}^{3\ell})$ adjoint of $a_{\ell}^{\dagger}$.
@@
$\mathcal{H}$ $=\sum_{i,j=1}^{d}{h}_{ij}a_{i}^{\dagger}a_{j}+{\frac{1}{2}}\sum_{i,j,k,\ell=1}^{d}{V}_{ijk\ell}a_{i}^{\dagger}a_{j}^{\dagger}a_{\ell}a_{k},$
where, for all $i,j,k,\ell\in\{1,\ldots,d\}$, it holds that
${h}_{ij}:=$ ${\frac{1}{2}}\int_{\mathbb{R}^{3}}\nabla\phi_{i}(\mathbb{x})\cdot\nabla\phi_{j}(\mathbb{x})\;d\mathbb{x}-\int_{\mathbb{R}^{3}}\sum_{\alpha=1}^{M}\frac{Z_{\alpha}}{|R_{\alpha}-\mathbb{x}|}\phi_{i}(\mathbb{x})\phi_{j}(\mathbb{x})\;d\mathbb{x},$
$\quad{V}_{ijk\ell}:=$ $\int_{\mathbb{R}^{3}}\int_{\mathbb{R}^{3}}\frac{\phi_{i}(\mathbb{x})\phi_{j}(\mathbb{y})\phi_{k}(\mathbb{x})\phi_{\ell}(\mathbb{y})}{|\mathbb{x}-\mathbb{y}|}\;d\mathbb{x}d\mathbb{y}.$
The ground-state problem for a molecular system composed precisely of $N$ electrons consists of finding the lowest eigenvalue of the second quantized Hamiltonian $\mathcal{H}$, subject to the constraint that the corresponding eigenvector describes an $N$-electron system. This constraint can be imposed by requiring that the sought energy minimizer be an eigenfunction of the so-called particle number operator $\mathcal{N}=\sum_{j=1}^{d}a_{j}^{\dagger}a_{j}\colon\mathcal{F}_{d}\rightarrow\mathcal{F}_{d}$ with eigenvalue precisely $N$. In other words, we seek
(21) $\mathcal{E}_{\rm GS}=\min_{0\neq\Phi\in\mathcal{F}_{d}}\frac{\langle\Phi,\mathcal{H}\Phi\rangle_{\mathcal{F}_{d}}}{\langle\Phi,\Phi\rangle_{\mathcal{F}_{d}}}\quad\text{subject to the constraint }\mathcal{N}\Phi:=\sum_{j=1}^{d}a_{j}^{\dagger}a_{j}\Phi=N\Phi.$
@@
Consequently, the minimization problem (21), which is a priori posed on the Fock space $\mathcal{F}_{d}$ can be reformulated over the space $\mathbb{R}^{2^{d}}=\bigotimes_{j=1}^{d}\mathbb{R}^{2}$ of order-$d$ tensors:
@@
The minimization problem (22) can now be approximated using the tensor train format and alternating minimization algorithms described in detail in Sections 2 and 3. Note, however, that these sections did not include any discussion on the imposition of the particle number constraint. Let us, therefore, briefly describe how this additional constraint can be dealt with in the tensor train format.
It has been shown by Bachmayr, Götte, and Pfeffer [1] that eigenfunctions of the particle number operator $\iota\circ\mathcal{N}\circ\iota^{-1}\rightarrow\mathbb{R}^{2^{d}}\times\mathbb{R}^{2^{d}}$ are characterized by a block-sparse structure in their tensor-train decomposition. Therefore, to approximate the minimization problem (22) on a tensor manifold $\mathcal{T}_{\mathbb{r}},\mathbb{r}\in\mathbb{N}^{d-1}$ using, e.g., classical one-site DMRG, we first initialize the algorithm with some $\mathbb{U}^{(0)}\in\mathcal{U}_{\mathbb{r}}$ that possess the requisite block-sparse structure. This guarantees that $\tau(\mathbb{U}^{(0)})$ is an eigenfunction of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$. Next, assume that we are performing micro-iteration $k+1$ of the one-site DMRG algorithm, and that the previous iterate $\mathbb{U}^{(k)}\in\mathcal{U}_{\mathbb{r}}$ is $k+1$-orthogonal and possesses the requisite block-sparse structure. As an alternative to Lemma 10, we can characterize the $k+1$ DMRG micro-iterate $\mathbb{U}^{(k+1)}\in\mathcal{U}_{\mathbb{r}}$ through projection operators on the tangent space $\mathbb{T}_{\tau(\mathbb{U}^{(k)})}\mathcal{T}_{\mathbb{r}}$.
To this end, for each $j\in\{1,\ldots,d\}$, let us denote by $\mathbb{U}^{(k)}_{\leq j}\in\mathbb{R}^{n_{1}\ldots n_{i}\times r_{i}}$ and $\mathbb{U}^{(k)}_{\geq j}\in\mathbb{R}^{r_{j-1}\times n_{j}\ldots n_{d}}$ the matrices obtained by contracting and reshaping the first $j$ TT cores and the last dj+1d-j+1 TT cores, respectively, of $\mathbb{U}^{(k)}$. We then introduce, for each $j\in\{1,\ldots,d\}$ the projection operators $\mathbb{P}_{\leq j},\mathbb{P}_{\geq j},\mathbb{P}_{j}\colon\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{T}_{\tau(\mathbb{U}^{(k)})}\mathcal{T}_{\mathbb{r}}$ as
@@
Here, $\text{\rm Ten}_{j}$ is a reshape operation that takes a matrix in $\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{d}}$ and reshapes it into a tensorial element of $\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$, and we adopt the convention that $\mathbb{P}_{\leq 0}=\mathbb{P}_{\geq d+1}=1$. Recalling Lemma 10, it is now possible to show (see, e.g., [1]) that
$\tau(\mathbb{U}^{(k+1)})\text{ is the lowest eigenfunction of }\mathbb{P}_{j}\circ\iota\circ\mathcal{H}\circ\iota^{-1}\circ\mathbb{P}_{j}^{*}.$
@@
Finally, we turn to the question of particle-number preservation in the A2DMRG Algorithm 2. As in the classical DMRG algorithms, we begin with an initialization $\mathbb{U}^{(0)}\in\mathcal{U}_{\mathbb{r}}$ that possesses the requisite block-sparse structure which guarantees that $\tau(\mathbb{U}^{(0)})$ is an eigenfunction of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$. We then perform $d$ (resp. d1d-1 in the two-site case) parallel DMRG micro-iterations resulting in updated tensors $\{\tau(\mathbb{V}^{(j)})\}_{j=1}^{d}$ (resp. $\{\tau(\mathbb{W}^{(j,j+1)})\}_{j=1}^{d-1}$). As argued above, each DMRG micro-iteration with an appropriately orthogonalized and particle-preserving initialization results in a particle-preserving tensor. Therefore, all tensors $\{\tau(\mathbb{V}^{(j)})\}_{j=1}^{d}$ (resp. $\{\tau(\mathbb{W}^{(j,j+1)})\}_{j=1}^{d-1}$) are eigenfunctions of $\iota\circ\mathcal{N}\circ\iota^{-1}$ with eigenvalue $N$ and so are linear combinations of such tensors. It therefore suffices to consider the effect of the compression step (Step 4 in Algorithm 2) which projects a linear combination $c^{*}_{0}\tau(\mathbb{U}^{(0)})+\sum_{j=1}^{d}c_{j}^{*}\tau(\mathbb{V}^{(j)}),~\{c_{j}^{*}\}_{j=0}^{d}\subset\mathbb{R}$ (resp. $d_{0}^{*}\tau(\mathbb{U}^{(0)})+\sum_{j=1}^{d-1}d_{j}^{*}\tau(\mathbb{W}^{(j,j+1)}),~\{d_{j}^{*}\}_{j=0}^{d-1}\subset\mathbb{R}$) onto some low-rank tensor manifold $\mathcal{T}_{\widetilde{\mathbb{r}}}$. Such a compression step can be performed using either the particle-preserving version of the TT-rounding algorithm [1, Algorithm 3], or more-generally, by solving a best-approximation problem using standard tensor train algorithms (such as alternating least squares), which are known to be particle-preserving [1, Section 6].
@@ 4. Numerical Experiments
For the purpose of the subsequent numerical study, we consider the tensor formulation of the ground state energy minimization problem (22) in the case of several so-called strongly correlated molecular systems. This includes linear chains of hydrogen atoms, i.e., molecular systems of the form H2p for $p\in\{3,4,5,6\}$ as well as the C2 and N2 dimers. The bond lengths in each Hydrogen chain and the C2 dimer are set to one Angstrom while the bond length for the N2 dimer is set to $1.0976$ Angstrom. For each molecular system, the $L^{2}(\mathbb{R}^{3})$-orthonormal single-particle basis $\{\phi_{j}\}_{j=1}^{d}$ is obtained by diagonalizing the Hartree-Fock (or Mean-Field) operator in the minimal STO-3G atomic basis. This implies, in particular, the relation $d=4p$ for each $p\in\{3,\ldots,6\}$ and $d=20$ for C2 and N2 so that our numerical experiments involve tensors of minimal order $d=12$ in the case of H6 and maximal order $d=24$ in the case of H12. The input data, i.e., the second quantized electronic Hamiltonian, and the reference minimal energies are obtained from the PySCF Python module for quantum chemistry [42] while the DMRG simulations themselves are run using an in-house Julia module developed and described in [2].
Figure 1. Convergence plot of the classical two-site DMRG and two-site A2DMRG algorithms for a linear, stretched H6 and H8 chain. The tolerances for the eigensolvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive half-sweeps/global iterations was less than $10^{-6}$.
@@
Figure 2. Convergence plot of the classical two-site DMRG and two-site A2DMRG algorithms for a linear, stretched H10 and H12 chain. The tolerances for the eigenvalue solvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive half-sweeps/global iterations was smaller than $10^{-6}$.
Figures 1-3 display the resulting error plots as a function of either the number of global iterations or the (dimensionless) cost per processor. The term global iteration refers to one half-sweep of the classical DMRG algorithm or one iteration of the A2DMRG Algorithm 2, both of which consists of solving exactly d1d-1 local minimization problems. The cost per processor is computed by evaluating, through the course of each global iteration, the number and flop-count of tensor contractions, matrix-vector products, and linear algebra operations such as QR and singular value decompositions. In the case of the classical DMRG algorithm, which is sequential in nature, we assume that all operations are performed on a single processor. In the case of the A2DMRG Algorithm 2, we assume that Steps 2 and 3, i.e., the DMRG micro-iterations and the solution of the second-level minimization problems are computed as much as possible in parallel. More precisely, we assume that each of the (d1)(d-1) DMRG micro-iterations in Step 2 of Algorithm 2 is performed on an independent processor, and each entry of the $d\times d$ symmetric coarse matrices appearing in the second-level minimization problem (Recall Equation 18) is computed on an independent processor. The cost per processor is then computed by taking the processor with the highest cost for Steps 2 and 3, and adding to it the cost of Steps 1 and 4 of Algorithm 2, the latter two steps being sequential in nature.
Figure 3. Convergence plot of the classical DMRG and A2DMRG algorithms for the C2 and N2 dimers. The tolerances for the eigenvalue solvers and the truncated SVD were set to $10^{-6}$ and the algorithms were run until the relative energy difference between successive half-sweeps/global iterations was smaller than $10^{-6}$.
@@
E The classical two-site DMRG algorithm.
F The two-site A2DMRG algorithm.
@@
A number of observations can now be made. First we note that the A2DMRG Algorithm 2 demonstrates slower convergence than the classical DMRG algorithm 1 as a function of the global iterations. Although, a rigorous quantification of the relative convergence rates of these two algorithms seems out of reach at the moment (indeed, even for the one-site DMRG algorithm, the local convergence analysis in [38] does not yield convergence rates), we feel that there is an intuitive explanation for the observed behavior: First, due to the sequential nature of the classical DMRG micro-iterations, the TT ranks of the ansatz grow very quickly and reach the maximal rank parameter $\mathbb{r}$ within two half-weeps. On the other hand, the A2DMRG algorithm requires several global iterations before a maximal rank parameter is achieved thereby slowing the convergence rate. Second, it is well-known in the domain decomposition literature that multiplicative Schwarz methods (which can be seen as generalizations of the Gauss-Seidel method) converge, in general, faster than additive Schwarz methods (which can be seen as generalizations of the Jacobi method). Intuitively, therefore, the classical DMRG algorithm which can be viewed as a non-linear, overlapping multiplicative Schwarz method should also converge faster than the A2DMRG Algorithm 2 which can be interpreted as a non-linear, overlapping additive Schwarz method. Interestingly, such a domain-decomposition-based interpretation of DMRG has– to our knowledge– not been studied in the literature.
Second, in spite of the previous observation on its relative convergence rate, we nevertheless observe that the A2DMRG Algorithm 2 exhibits a significant parallel speedup in terms of the cost per processor. The exact speed-up varies depending on the choice of ranks and molecule (the latter determines the order $d$ of the tensor), but ranges from a factor two to a factor six. Part of the reason for this speed-up (which is lower than the theoretical optimal speed-up of factor $d/4$) is that the classical and additive DMRG algorithms seem to require a different number of Lanczos iterations for the solution of the micro-iterations. To explore this, we plot in Figure 4, the number of Lanczos iterations required for the DMRG micro-iterations over the course of the DMRG half-sweeps/global iterations for several of these molecular systems. These results indicate that, in general, only a minority of local problems in the classical DMRG algorithm require more than a single Lanczos iteration. This is not the case for the A2DMRG algorithm, which typically requires several Lanczos iterations for each micro-iteration, thereby producing a non-optimal speed-up. However, we emphasize that the number of Lanczos iterations for the A2DMRG algorithm displays minimal growth as a function of the maximal TT ranks $\mathbb{r}$ and the order $d$ of the tensor. The corresponding results for H8, H10 and the N2 dimer are similar and are therefore not displayed for brevity.